# Promoted current-induced spin polarization in inversion symmetry broken topological insulator thin films

### v1

@@ Abstract
We theoretically investigate current-induced spin polarization in disordered topological insulator thin films with broken inversion symmetry under an applied in-plane electric field. Utilizing the Kubo formalism within the self-consistent Born approximation—and incorporating vertex corrections to account for multiple scattering events—we analyze how disorder, chemical potential, the electrostatic potential difference between the top and bottom surfaces, and momentum-dependent hybridization affect the spin susceptibility. Our results reveal that the spin susceptibility exhibits nonzero values within a finite range around a zero gap, and this range broadens as the chemical potential increases. A higher hybridization strength induces asymmetry in the spin response, and a stronger potential difference, breaking inversion symmetry, significantly enhances polarization—a trend attributable to band inversion that is further refined by vertex corrections. These findings provide a theoretical framework for tuning spin-charge conversion in topological thin films, with implications for spintronic device applications.
pacs:
I Introduction
@@
As already mentioned above, the main reason for the existence of TIs is the spin-orbit coupling [26], which is a combination of spin- and orbital-freedom degrees resulting in various attractive effects [27]. One of the most interesting phenomena related to spin-orbit coupling is transverse spin flows [28]. When an electric field is applied to a system in the presence of spin-orbit coupling, a transverse spin current appears, known as the spin Hall effect (SHE) [29]. In SHE, a net spin current is produced perpendicular to the electric field, leading to the accumulation of spins at the edges of the sample [30, 31]. In addition, this applied electric field can also cause the spin of the charge carriers to polarize when the carriers flow through a nonmagnetic medium [32]. This phenomenon, known as the Edelstein effect [33] or current-induced spin polarization [34], arises from the asymmetric relaxation of different spin states [35, 36]. In practice, it can emerge through the locking of the spin to its momentum. Note that the current-induced spin polarization is basically different from the SHE. In the case of current-induced spin polarization, a homogeneous spin polarization of electrons appears throughout the sample [32, 37].
@@
Current-induced spin polarization has been observed in various environments experimentally, including semiconductors [38, 39, 40, 41, 42, 43, 44, 45], van der Waals heterostructures [46, 47] and heavy metal thin films [48, 49, 50]. It has been studied theoretically in a wide range of systems, ranging from two-dimensional electron gases [51, 52, 53, 54] to three-dimensional (bulk) materials [55, 56, 57, 58]. It has been shown that structure-property determines configurations of charge-to-spin conversion [57]. For example, Dyrdał et al. has investigated the current-induced spin polarization in graphene in the presence of the Rashba spin–orbit interaction, demonstrating that the spin polarization is perpendicular to the electric field and confined to the graphene plane [59]. More recently, this work has been extended, revealing that the direction of spin polarization can be tuned by the anisotropy of the Rashba interaction [60].
Spin manipulation [61, 62], modulated by electrical means, plays an essential role in the advancement of spintronics [63], particularly with the integration of novel materials such as TI [64] due to the high spin-orbit coupling in these materials [65, 66]. Recent research has extensively investigated current-induced spin polarization in 3D TIs [67, 68, 69, 70, 71, 72], even at room temperature [73, 74], paving the way for the realization of the TI spin transistor [75]. Current-induced spin polarization has been shown to be observed in TIs using a circularly polarized photoconductive differential current [76] and ferromagnetic resonance [77]. Furthermore, a comparison of current-induced spin polarization in the TI Bi2Se3 and InAs Rashba states has been made, indicating a dominant contribution to current-induced spin polarization in TI [78]. Also, the combined effect of current-induced spin polarization and scattering processes can result in bilinear magnetoresistance in TIs due to the inhomogeneities of spin-momentum locking [79]. Hybrid structures of TI thin films and magnetic insulators have been shown to exhibit an anisotropic Edelstein effect [80]. After all, considering the features of a TI thin film, its current-induced spin polarization is relatively less studied [81], and it really deserves further research.
@@
Despite numerous studies on spin-charge interconversion in topological materials, a comprehensive understanding of current-induced spin polarization in thin films with broken inversion symmetry remains incomplete—particularly in the presence of both disorder and hybridization. Here, we address this by developing a theoretical model incorporating self-consistent impurity effects and vertex corrections. Our approach reveals how structural asymmetry and momentum-dependent coupling modulate the spin response, with implications for spintronics.
II Model
@@
Figure 1: (Color online) Schematic diagram of a TI thin film subjected to a parallel applied electric field $E$ in the $x$-direction and an electrostatic potential $V$ between the top and bottom surfaces. Also, $\Delta_{0}$ and $\Delta_{1}$ are hybridization parameters between the two surface states. The homogeneous spin polarization of electrons across the sample, resulting from the current-induced spin polarization effect, is depicted in blue.
@@
We consider a TI with thickness less than 6 quintuple layers to study current-induced spin polarization in the TI thin film as indicated schematically in Fig. 1. The TI thin film is assumed to be under an electric field, parallel to the top and bottom surfaces in a given direction such as $x$-direction. In addition, an electrostatic potential is applied perpendicularly to the surfaces in the z-direction breaking inversion symmetry. In fact, this system is similar to a bilayer in which the top and bottom surface states can hybridize. Hence, the corresponding Hamiltonian can be written as [18, 82, 83, 84];
@@
where $\sigma$ and $\tau$ are the Pauli matrices in spin and surface spaces, $k$ is in-plane momentum, and $v=\hbar\upsilon_{f}$ with $\upsilon_{f}$ being the Fermi velocity. $\Delta$ is the element of the hybridization matrix of the top and bottom surface states that is usually shown as $\Delta=\Delta_{0}-\Delta_{1}k^{2}$. Here, $\Delta_{0}$ denotes the zero-momentum hybridization energy, while $\Delta_{1}$ characterizes the momentum-dependent correction; these parameters determine the gap structure observed in the surface states. Also, $V$ is the difference in electrostatic potential between the top and bottom surfaces. The parameters $V$, $\Delta_{0}$, and $\Delta_{1}$ determine whether the system is in a trivial or nontrivial state. When $\Delta_{0}\Delta_{1}>0$ and $V<V_{c}$, with $V_{c}=v\sqrt{|{\Delta_{0}/\Delta_{1}}|}$, being a critical potential difference, the system is topologically nontrivial. But when the potential different increases, $V>V_{c}$, the system faces a transition from topological to normal insulator [84].
@@
At $V=V_{c}$, the band gap closes and the Dirac point splits into two gapless Dirac points separated horizontally with finite momentum ±$\pm k_{D}=\pm V_{c}/v$ giving rise to asymmetric band dispersion. The band touching points are located symmetrically around $k=0$, as shown in Fig. 2(b). The disappearance of the gap indicates that the hybridization and potential difference are exactly balanced, setting the stage for the transition between a topological insulator to a trivial insulator or vice versa. At this point, the band inversion is about to reverse, and the surface state symmetry is delicately balanced, causing strong sensitivity to small perturbations. Spin polarization can undergo abrupt changes here as a result of the degeneracy of spin-momentum locked states.
@@ III Effect of impurities
$\Sigma^{\pm}=<\tilde{V}G^{\pm}\tilde{V}>=n_{i}V_{0}^{2}\int\frac{d^{2}k}{(2\pi)^{2}}G^{\pm}.$ (9)
Based on the above equation and according to Eq.(7), the nonzero components of the self-energy can be found as $\Sigma_{00}^{\pm}=\Sigma_{xx}^{\pm}$, $\Sigma_{yy}^{\pm}=\Sigma_{zz}^{\pm}$, and $\Sigma_{0y}^{\pm}=\Sigma_{y0}^{\pm}=\Sigma_{xz}^{\pm}=\Sigma_{zx}^{\pm}$ with
@@
Equations (7)–(13) describe the impurity-averaged Green’s function and self-energy corrections in detail. First, we adopt the self-consistent Born approximation, which assumes weak impurity scattering and allows us to treat disorder perturbatively; second, these equations capture both the energy renormalization and the disorder-induced broadening of the electronic states; and finally, they reveal that hybridization between the top and bottom surface states, in conjunction with impurity scattering, plays a pivotal role in modifying the band structure and consequently impacts the spin-polarization response in the TI thin film. Thus, for clarity, we decompose the self-energy as $\Sigma^{\pm}=\Sigma^{\prime}\mp i\Gamma$ where $\Sigma^{\prime}$ and $\Gamma$ are the real and imaginary parts of the self-energy. Within the self-consistent treatment, we evaluated the components of the self-energy Eqs. (10-12).
Figure 3 presents the real and imaginary parts of the self-energy components $\Sigma_{00}^{-}$, $\Sigma_{0y}^{-}$ and $\Sigma_{zz}^{-}$ for a thin film of TI. Panel (a) shows the real parts of self-energy terms. The real parts of the self-energy components represent the energy shifts resulting from impurity scattering and hybridization effects. Variations among $\Sigma_{00}^{{}^{\prime}}$, $\Sigma_{0y}^{{}^{\prime}}$ and $\Sigma_{zz}^{{}^{\prime}}$ indicate that impurities renormalize the quasi-particle energies in a nonuniform (anisotropic) way. $\Sigma_{0y}^{{}^{\prime}}$ behaves differently than $\Sigma_{zz}^{{}^{\prime}}$ suggesting that the in-plane and out-of-plane scattering channels are affected differently by the disorder.
Panel (b) displays the imaginary parts of the self-energies. The imaginary parts of the self-energy components are directly related to the scattering rate and, hence, the finite lifetime of the electronic states. A higher magnitude in these components indicates stronger disorder-induced broadening. The distinct behavior of the different components implies that not all scattering processes equally damp the electronic states. In particular, $\Gamma_{0y}$ and $\Gamma_{zz}$ show different trends. It suggests that the decay rates (or scattering lifetimes) along different directions are not identical, reinforcing the notion of anisotropic scattering. As a result, the real and imaginary parts of the self-energy components reveal channel-dependent energy shifts and scattering rates. Notably, the anisotropy in these components, particularly at finite inversion-breaking potentials, implies direction-dependent relaxation dynamics. This could manifest as anisotropic spin conductivity and relaxation times, which are critical in practical TI-based spintronic systems.
@@
IV Current induced spin polarization
@@
[…]
where $f(E)$ is the Fermi distribution function, and $v_{\beta}=(\partial H/\partial k_{\beta})/\hbar$ is the velocity matrix along the $\beta$-th axis, with $\beta=x,y$. $Tr$ stands for trace over spin space. Taking into account vertex corrections [33, 88] in the ladder approximation, Eq. (15) for $\chi_{\alpha\beta}$ at zero temperature can be recast as
@@
[…]

### v2

@@ Abstract
We theoretically investigate current-induced spin polarization in disordered topological insulator thin films with broken inversion symmetry under an applied in-plane electric field. Utilizing the Kubo formalism within the self-consistent Born approximation and incorporating vertex corrections to account for multiple scattering events, we analyze how disorder, chemical potential, the electrostatic potential difference between the top and bottom surfaces, and momentum-dependent hybridization affect the spin susceptibility. Our results reveal that the spin susceptibility exhibits nonzero values within a finite range around a zero gap, and this range broadens as the chemical potential increases. A higher hybridization strength induces asymmetry in the spin response. A stronger potential difference, breaking inversion symmetry, significantly enhances polarization. This enhancement is a trend attributable to band inversion and is further refined by vertex corrections. These findings provide a theoretical framework for tuning spin-charge conversion in topological thin films, with implications for spintronic device applications.
keywords:
Spin Polarization , Topological Insulator , Thin Film
1 Introduction
@@
As already mentioned above, the main reason for the existence of TIs is the spin-orbit coupling [26], which is a combination of spin- and orbital-freedom degrees resulting in various attractive effects [27]. One of the most interesting phenomena related to spin-orbit coupling is transverse spin flows [28]. When an electric field is applied to a system in the presence of spin-orbit coupling, a transverse spin current appears, known as the spin Hall effect [29]. In the spin Hall effect, a net spin current is produced perpendicular to the electric field, leading to the accumulation of spins at the edges of the sample [30, 31]. In addition, this applied electric field can also cause the spin of the charge carriers to polarize when the carriers flow through a nonmagnetic medium [32]. This phenomenon, known as the Edelstein effect [33] or current-induced spin polarization (CISP) [34], arises from the asymmetric relaxation of different spin states [35, 36]. In practice, it can emerge through the locking of the spin to its momentum. Note that the CISP is basically different from the SHE. In the case of CISP, a homogeneous spin polarization of electrons appears throughout the sample [32, 37].
@@
CISP has been observed in various environments experimentally, including semiconductors [38, 39, 40, 41, 42, 43, 44, 45], van der Waals heterostructures [46, 47] and heavy metal thin films [48, 49, 50]. It has been studied theoretically in a wide range of systems, ranging from two-dimensional electron gases [51, 52, 53, 54] to three-dimensional (bulk) materials [55, 56, 57, 58, 59]. It has been shown that structure-property determines configurations of charge-to-spin conversion [58]. For example, Dyrdał et al. has investigated the CISP in graphene in the presence of the Rashba spin–orbit interaction, demonstrating that the spin polarization is perpendicular to the electric field and confined to the graphene plane [60]. More recently, this work has been extended, revealing that the direction of spin polarization can be tuned by the anisotropy of the Rashba interaction [61]. Spin manipulation [62, 63], modulated by electrical means, plays an essential role in the advancement of spintronics [64], particularly with the integration of novel materials such as TI [65] due to the high spin-orbit coupling in these materials [66, 67].
CISP in three-dimensional TIs has been studied both theoretically and experimentally. Recent works revealed the underling mechanisms of CISP in TI surface states [68]. However, subsequent experiments demonstrated its detection on various material platforms and device architectures [69, 70, 71]. Room-temperature nonlocal spin–charge interconversion in Rashba-type systems and TI-based heterostructures was reported [72, 73]. Gate-controlled sign reversal of spin polarization has been realized in TI spin transistors [74]. Several complementary detection schemes have been developed, including photoconductive differential current measurements [75], ferromagnetic resonance [76], transport in tunnel-coupled interfaces [70, 71], and opposite spin polarizations in bulk-insulating and bulk-metallic TI flakes [77]. Furthermore, a comparison of CISP has been made in the TI Bi2Se3, InAs Rashba states [78], and modulated Dirac surface states [79], indicating a dominant contribution of TI surface states to CISP. Furthermore, Hwang et al. detected electrically generated spin polarization in bulk-insulating Bi1.5Sb0.5Te1.7Se1.3 [80].
Burkov et al. formulated diffusion equations for spin–charge coupled transport on TI surfaces. They showed that strong spin–momentum locking gives rise to a novel magnetoresistance effect, with the response tunable by gate voltage and relevant for spin-transistor applications [81]. Through a linear-response density-matrix formalism, it has been shown that the Edelstein effect in TI surface states remains robust in the presence of electron–electron interactions and disorder, although both conductivity and spin polarization are renormalized [82]. Furthermore, a comprehensive description of spin–charge interconversion in TI thin films has been provided, clarifying how structural asymmetry and surface hybridization can be exploited to improve spin–orbit torques and optimize device performance [83]. In contrast, Li et al. argued within a Boltzmann transport approach that the net ensemble spin polarization induced by current is intrinsically small, cautioning against over-interpretation of experimental signals [84]. Using a semiclassical treatment, it has been shown that combining impurity scattering with spin-momentum locking of the Dirac surface states, the CISP is independent of chemical potential and temperature [85]. Moreover, TI/ferromagnet heterostructures have been shown to exhibit anisotropic Edelstein effects and tunable spin–orbit torques [86]. Spin-momentum locking inhomogeneities have also been identified as a source of bilinear magnetoresistance, providing new insight into nonlinear spin–charge transport [87]. This provides an efficient charge-to-spin conversion in TIs [88] even in the presence of the anisotropic Fermi contours [89].
@@
In contrast, in ultrathin films, hybridization between top and bottom surfaces and inversion symmetry breaking introduce qualitatively new features in the spin response. Considering the features of a TI thin film, its CISP is relatively less studied [90], and it really deserves further research. Here, we address this by developing a theoretical model that incorporates self-consistent impurity effects and vertex corrections. Our approach reveals how structural asymmetry and momentum-dependent coupling modulate the spin response, with implications for spintronics.
2 Model
@@
Figure 1: (Color online) Schematic diagram of a TI thin film subjected to a parallel applied electric field $E$ in the $x$-direction and an electrostatic potential $V$ between the top and bottom surfaces. Also, $\Delta_{0}$ and $\Delta_{1}$ are hybridization parameters between the two surface states. The homogeneous spin polarization of electrons across the sample, resulting from the CISP effect, is depicted in blue.
@@
We consider a TI with thickness less than 6 quintuple layers to study CISP in the TI thin film as indicated schematically in Fig. 1. The TI thin film is assumed to be under an electric field, parallel to the top and bottom surfaces in a given direction such as $x$-direction. In addition, an electrostatic potential is applied perpendicularly to the surfaces in the z-direction breaking inversion symmetry. In fact, this system is similar to a bilayer in which the top and bottom surface states can hybridize. Hence, the corresponding Hamiltonian can be written as [18, 91, 92, 93];
@@
where $\sigma$ and $\tau$ are the Pauli matrices in spin and surface spaces, $k$ is in-plane momentum, and $v=\hbar\upsilon_{f}$ with $\upsilon_{f}$ being the Fermi velocity. $\Delta$ is the element of the hybridization matrix of the top and bottom surface states that is usually shown as $\Delta=\Delta_{0}-\Delta_{1}k^{2}$. The hybridization parameters $\Delta_{0}$ and $\Delta_{1}$ describe the tunneling between the top and bottom surfaces of the TI thin film. $\Delta_{0}$ denote the gap at zero momentum, while $\Delta_{1}$ characterizes how the gap changes with momentum. Also, $V$ is the difference in electrostatic potential between the top and bottom surfaces. The electrostatic potential difference $V$ breaks inversion symmetry, changing the states of the two surfaces. The parameters $V$, $\Delta_{0}$, and $\Delta_{1}$ determine whether the system is in a trivial or nontrivial state. When $\Delta_{0}\Delta_{1}>0$ and $V<V_{c}$, with $V_{c}=v\sqrt{|{\Delta_{0}/\Delta_{1}}|}$, being a critical potential difference, the system is topologically nontrivial, corresponds to a quantum spin Hall phase of the effective 2D system. But when the potential different increases, $V>V_{c}$, the system faces a transition from topological to normal insulator [93].
It should be noted that the topology discussed here refers to the effective two-dimensional thin-film system formed by the hybridized top and bottom surface states, and not the original three-dimensional bulk topology of the parent topological insulator. The bulk is assumed to be a strong 3D TI, ensuring the existence of gapless surface states in the absence of hybridization. In the thin-film geometry, these states couple through tunneling, resulting in a massive Dirac Hamiltonian with the hybridization matrix term $\Delta$. Following the analysis of [93, 94], $\mathbb{Z}_{2}$ invariant of this effective 2D model can be computed using the Pfaffian method.
@@
At $V=V_{c}$, the band gap closes and the Dirac point splits into two gapless Dirac points separated horizontally with finite momentum $\pm k_{D}=\pm V_{c}/v$ giving rise to asymmetric band dispersion. The band touching points are located symmetrically around $k=0$, as shown in Fig. 2(b). The disappearance of the gap indicates that the hybridization and potential difference are exactly balanced, setting the stage for the transition between a topological insulator to a trivial insulator or vice versa. At this point, the band inversion is about to reverse, and the surface state symmetry is delicately balanced, causing strong sensitivity to small perturbations. Spin polarization can undergo abrupt changes here as a result of the degeneracy of spin-momentum locked states.
@@ 3 Effect of impurities
$\Sigma^{\pm}=\langle\tilde{V}G^{\pm}\tilde{V}\rangle=n_{i}V_{0}^{2}\int\frac{d^{2}k}{(2\pi)^{2}}G^{\pm}.$ (9)
The Green function on the right-hand side of Eq. (9) has a matrix structure originating from the tensor product space of the layer and the spin degrees of freedom. So, the self-energy has the $\times4$ matrix structure as
$\Sigma^{\pm}=\begin{bmatrix}\Sigma^{\pm}_{00}&\Sigma^{\pm}_{0x}&\Sigma^{\pm}_{0y}&\Sigma^{\pm}_{0z}\\ \Sigma^{\pm}_{x0}&\Sigma^{\pm}_{xx}&\Sigma^{\pm}_{xy}&\Sigma^{\pm}_{xz}\\ \Sigma^{\pm}_{y0}&\Sigma^{\pm}_{yx}&\Sigma^{\pm}_{yy}&\Sigma^{\pm}_{yz}\\ \Sigma^{\pm}_{z0}&\Sigma^{\pm}_{zx}&\Sigma^{\pm}_{zy}&\Sigma^{\pm}_{zz}\end{bmatrix}$, (10)
where $\Sigma^{\pm}_{ij}$ with $i,j=0,x,y,z$ are the components of self-energy. Note that, in our definition, the first index refers to the layer component in $\tau$-space, and the second to the spin component in $\sigma$-space, not in the crystal axes. Based on the above equation and according to Eq. (7), the nonzero components of the self-energy can be found as $\Sigma_{00}^{\pm}=\Sigma_{xx}^{\pm}$, $\Sigma_{yy}^{\pm}=\Sigma_{zz}^{\pm}$, and $\Sigma_{0y}^{\pm}=\Sigma_{y0}^{\pm}=\Sigma_{xz}^{\pm}=\Sigma_{zx}^{\pm}$ with
@@
Equations (7)–(14) describe the impurity-averaged Green’s function and self-energy corrections in detail. First, we adopt the self-consistent Born approximation, which assumes weak impurity scattering and allows us to treat disorder perturbatively; second, these equations capture both the energy renormalization and the disorder-induced broadening of the electronic states; and finally, they reveal that hybridization between the top and bottom surface states, in conjunction with impurity scattering, plays a pivotal role in modifying the band structure and consequently impacts the spin-polarization response in the TI thin film. Thus, for clarity, we decompose the self-energy as $\Sigma^{\pm}=\Sigma^{\prime}\mp i\Gamma$ where $\Sigma^{\prime}$ and $\Gamma$ are the real and imaginary parts of the self-energy. Within the self-consistent treatment, we evaluated the components of the self-energy Eqs. (11-13).
@@
Figure 3 presents the real and imaginary parts of the self-energy components $\Sigma_{00}^{-}$, $\Sigma_{0y}^{-}$ and $\Sigma_{zz}^{-}$ for a thin film of TI under inversion symmetry breaking. Panel (a) shows the real parts of self-energy terms. The real parts of the self-energy components represent the energy shifts resulting from impurity scattering and hybridization effects. Variations among $\Sigma_{00}^{{}^{\prime}}$, $\Sigma_{0y}^{{}^{\prime}}$ and $\Sigma_{zz}^{{}^{\prime}}$ indicate that impurities renormalize different layer–spin channels in a nonuniform way. In particular, the difference between $\Sigma_{0y}^{{}^{\prime}}$ and $\Sigma_{zz}^{{}^{\prime}}$ suggests that spin channels coupled to $\sigma_{y}$ and $\sigma_{z}$ respond differently to disorder, even though the inversion-breaking potential is spin independent. Because, the potential modifies the states of the top and bottom surfaces. In the presence of spin–momentum locking, the surface-state spin texture depends on momentum and couples differently to $\sigma_{x}$ and $\sigma_{y}$ spin components. The momentum-dependent hybridization term $\Delta_{1}k^{2}$ enhances this difference, leading to unequal self-energy components.
Panel (b) displays the imaginary parts of the self-energies. The imaginary parts of the self-energy components are directly related to the scattering rate and, hence, the finite lifetime of the electronic states. A higher magnitude in these components indicates stronger disorder-induced broadening. The distinct behavior of the different components implies that not all scattering processes equally damp the electronic states. In particular, $\Gamma_{0y}$ and $\Gamma_{zz}$ show different behaviours. It suggests that different spin–layer channels experience different levels of disorder-induced damping. In other words, the lifetime broadening is channel dependent, not uniform across all components.
[…]
4 Current induced spin polarization
@@
$\chi_{\alpha\beta}=-\frac{e\hbar}{2\pi}\int dE\frac{\partial f(E)}{\partial E}Tr\langle\hat{S}_{\alpha}G^{+}_{0}(E)v_{\beta}G^{-}_{0}(E)\rangle_{c},$ (16)
where $f(E)$ is the Fermi distribution function, and $v_{\beta}=(\partial H/\partial k_{\beta})/\hbar$ is the velocity matrix along the $\beta$-th axis, with $\beta=x,y$. $Tr$ stands for trace over spin space.
In disordered systems, impurity scattering modifies the electrons respond to external fields. In our calculation, this effect appears as a renormalization of the velocity operator, called the vertex correction. Diagrammatically, this corresponds to adding an infinite series of “ladder” diagrams, where the velocity operator is repeatedly connected by impurity lines to pairs of Green’s function. The ladder structure ensures that multiple scattering events are taken into account self-consistently. In practice, this means that the bare velocity $v_{\beta}$ is replaced by a dressed velocity $\tilde{v}_{\beta}$ obtained from the Bethe–Salpeter equation. Taking into account vertex corrections [33, 98] in the ladder approximation, Eq. (16) for $\chi_{\alpha\beta}$ at zero temperature can be recast as
@@
$\chi_{yx}=C\int dk\frac{8kV\mu(V^{2}-\mu^{2}+k^{2}v^{2}+\Delta^{2})(V^{2}-\mu^{2}-k^{2}v^{2}+\Delta(\Delta_{0}+3\Delta_{1}k^{2}))}{\prod_{n=1}^{4}(\mu-E_{n}+i\eta)(\mu-E_{n}-i\eta)},$